GISBERT WINNEWISSER: AN APPRECIATION

ERIC HERBST, Departments of Physics, Chemistry, and Astronomy, The Ohio State University, Columbus OH.

The death of Gisbert Winnewisser in March of this year has robbed us of a peerless spectroscopist and molecular astronomer who made many significant contributions to both subjects. Moreover, Gisbert was a friend and mentor to many of us, who enriched our lives immeasuarably. In this brief appreciation, I will discuss some of his many accomplishments.